Compute and find a relation between the expressions

(8.1) ~u · (~v × w~ ), w~ · (~u × ~v), and ~v · (w~ × ~u)

Find an expression for :

(8.2) ~u · (~v × w~ ) and ~u × (~v × w~ )

(8.3) Find the side lengths and angles of the triangle with vertices the tips the vectors in Question 6 above.

(8.4) Use ~u and ~v from Question 6 to find the area of the parallelogram formed by ~u and ~v.

(8.5) Use ~u, ~v and w~ from Question 6 to find the volume of the parallelepiped with edges deter- mined by the three vectors ~u, ~v and w~ 


1
Expert's answer
2021-06-07T16:58:56-0400

(8.1)

The scalar triple product is unchanged under a circular shift of its three operands


"\vec u \cdot(\vec v \times \vec w)=\vec w \cdot(\vec u \times \vec v)=\vec v \cdot(\vec w \times \vec u)"

"\vec u=\lang-1, 1, 2\rang, \vec v=\lang2, -1, 0\rang, \vec w=\lang1, 1, 3\rang"

"\vec u \cdot(\vec v \times \vec w)=\begin{vmatrix}
 -1 & 1 & 2 \\
 2 & -1 & 0 \\
 1 & 1 & 3 \\
\end{vmatrix}"

"=-1\begin{vmatrix}
 -1 & 0 \\
 1 & 3
\end{vmatrix}-1\begin{vmatrix}
 2 & 0 \\
 1 & 3
\end{vmatrix}+2\begin{vmatrix}
 2 & -1 \\
 1 & 1
\end{vmatrix}"

"=3-6+6=3"

"\vec u \cdot(\vec v \times \vec w)=\vec w \cdot(\vec u \times \vec v)=\vec v \cdot(\vec w \times \vec u)=3"

(8.2)


"\vec u \cdot(\vec v \times \vec w)=3"

"\vec v \times \vec w=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 2 & -1 & 0 \\
 1 & 1 & 3 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 -1 & 0 \\
 1 & 3
\end{vmatrix}-\vec j\begin{vmatrix}
 2 & 0 \\
 1 & 3
\end{vmatrix}+\vec k\begin{vmatrix}
 2 & -1 \\
 1 & 1
\end{vmatrix}"

"=-3\vec i-6\vec j+3\vec k"


"\vec u \times(\vec v \times \vec w)=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 -1 & 1 & 2 \\
 -3 & -6 & 3 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & 2 \\
 -6 & 3
\end{vmatrix}-\vec j\begin{vmatrix}
 -1 & 2 \\
 -3 & 3
\end{vmatrix}+\vec k\begin{vmatrix}
 -1 & 1 \\
 -3 & -6
\end{vmatrix}"

"=15\vec i-3\vec j+9\vec k"


(8.3)

"\vec u=\lang-1, 1, 2\rang, \vec v=\lang2, -1, 0\rang, \vec w=\lang1, 1, 3\rang"


"\overrightarrow{AB}=\lang3, -2, -2\rang, \overrightarrow{BC}=\lang-1, 2, 3\rang, \overrightarrow{AC}=\lang2, 0, 1\rang"


"AB=\sqrt{(3)^2+(-2)^2+(-2)^2}=\sqrt{17}"

"BC=\sqrt{(-1)^2+(2)^2+(3)^2}=\sqrt{14}"


"AC=\sqrt{(2)^2+(1)^2+(0)^2}=\sqrt{5}"

"\overrightarrow{AB}\cdot \overrightarrow{AC}=3(2)-2(0)-2(1)=4"

"\overrightarrow{BA}\cdot \overrightarrow{BC}=-3(-1)+2(2)+2(3)=13"

"\overrightarrow{CA}\cdot \overrightarrow{CB}=-2(1)+0(-2)-1(-3)=1"

"\cos\alpha=\dfrac{-3}{\sqrt{6}\sqrt{5}}=>\alpha=180\degree-\cos^{-1}\dfrac{\sqrt{30}}{10}"

"\approx64.3\degree"

"\cos\beta=\dfrac{13}{\sqrt{17}\sqrt{14}}=>\beta=\cos^{-1}\dfrac{13\sqrt{238}}{238}"

"\approx32.6\degree"

"\cos\gamma=\dfrac{1}{\sqrt{5}\sqrt{14}}=>\gamma=\cos^{-1}\dfrac{\sqrt{70}}{70}"

"\approx83.1\degree"

(8.4)


"\vec u \times \vec v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 -1 & 1 & 2 \\
 2 & -1 & 0 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & 2 \\
 -1 & 0
\end{vmatrix}-\vec j\begin{vmatrix}
 -1 & 2 \\
 2 & 0
\end{vmatrix}+\vec k\begin{vmatrix}
 -1 & 1 \\
 2 & -1
\end{vmatrix}"

"=2\vec i+\vec j-\vec k"

"Area=\|\vec u \times \vec v\|=\sqrt{(2)^2+(1)^2+(-1)^2}=\sqrt{6}\ (units^2)"


(8.5)


"V=|\vec u \cdot(\vec v \times \vec w)|=\bigg|\begin{vmatrix}
 -1 & 1 & 2 \\
 2 & -1 & 0 \\
 1 & 1 & 3 \\
\end{vmatrix}\bigg|=3 (units^3\ )"